Circular No. 6860

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

AX J0051–73.3 = RX J0050.7–7316 K. Cook, Lawrence Livermore National Laboratory, reports that the

MACHO collaboration (cf. IAUC 6312) has detected an eclipsing variable star with peculiar properties within 2′′.0 of the position for AX J0051– 73.3 reported on IAUC 6854. Located at α = 0^{h}50^{m}44^{s}.7, δ = −73^{o}16′05′′ (equinox J2000.0), this is the nearest source in our database to the position of the x-ray source and is well within the reported 90-percent-confidence region. The lightcurve has mean V = 14.7 and V − R = +0.2, with a pronounced, long-term baseline brightness variation of 0.1 mag from 1993 June 18.7 UT to the present. A persistent contact-binary-like lightcurve with period 0.708 days and peak-to-peak amplitude of 0.10 mag is seen superposed on the much longer timescale variation. The MACHO Project data are acquired at the 1.5-m telescope at Mount Stromlo Observatory.

GRB 980329 J. M. Quashnock, D. E. Vanden Berk, D. M. Cole, A. R. Cooray, D. Q.

Lamb, F. J. Castander, K. Gloria, and D. Long, University of Chicago, on behalf of the Astrophysical Research Consortium, report: “Between Apr. 1.167 and 1.277 UT, we made near-infrared observations of a region of the sky that includes the entire BeppoSAX NFI error box for GRB 980329 (IAUC 6853, 6854), using the GRIM II instrument mounted on the ARC 3.5-m telescope at Apache Point Observatory. During Apr. 1.167–1.177, at α = 7^{h}02^{m}38^{s}.7, δ = +38^{o}50′27′′ (equinox 2000.0) — consistent with the position of the object of mag R = 20.6 noted on IAUC 6856 — we find an object of mag J = 17.7 ± 0.1; during Apr. 1.267–1.277 (∼ 2.5 hr later), we find no measurable change (less than ∼ 0.1 mag) in the J magnitude of the object. This object is ∼ 6′′.4 south and 3′′.0 west of the bright (R = 15.8) star reported on IAUC 6856, for which we measure J = 13.9 ± 0.1.”

SUPERNOVA 1998S IN NGC 3877 P. Garnavich, S. Jha, and R. Kirshner, Harvard-Smithsonian Center for

Astrophysics, report: “Spectra of SN 1998S by P. Berlind and M. Calkins with the 1.5-m Tillinghast telescope show that a deep, broad absorption feature (with a minimum observed at 628.6 nm) has developed in the alreadyunusual type-II spectrum. Assuming the same expansion velocity as He I 587.5-nm (4600 km/s), the feature has a rest wavelength of 636 nm. This likely corresponds to the Si II blend (634.7, 637.1 nm) often seen in type-I supernovae.”

1998 April 3  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Daniel W. E. Green